### v1

# Construction of Exponential Families from Statistical Manifolds

Abstract.
This paper addresses the fundamental question of constructing exponential families from statistical manifolds. We show that every compact statistical manifold admits a singular foliation whose leaves are Hessian manifolds. Moreover, any non-flat, compact, orientable 3-dimensional leaf appears as a quotient of an exponential family and has only odd Betti.
1. Introduction
Information geometry is a rapidly developing discipline that establishes a profound connection between differential geometry and statistical theory through the Fisher information metric (see [amari2000methods, amari2016information, lauritzen1987statistical, ay2017information]). Numerous researchers have explored the construction of meaningful distances between probability distributions, acknowledging its critical role in diverse areas including theoretical statistics, information theory, signal processing, quantitative finance, and machine learning. To solve this problem, Rao [rao1945information] introduced the geodesic distance derived from the Fisher information metric, now universally recognized as the Fisher-Rao distance (or Rao’s distance). The Fisher metric, also called the information metric, is generally positive semi-definite but can be degenerate. Cencov ([cencov2000statistical], p. 156) proved the uniqueness result for Fisher’s information metric on finite sample spaces. Subsequently, in [ay2015information], the authors extended Cencov’s result to infinite sample spaces. The Fisher information metric is invariant under parametrization on a sample space $\Xi$ and covariant under parametrization on a parametric space $\Theta$, see [calin2014geometric]. The Fisher-Rao distance is inherently designed to be invariant under diffeomorphisms of both the sample space $\Xi$ and the parameter space $\Theta$. The Fisher metric has been widely used in many other areas of geometry (see [hitchin2006geometry, fujiwara2020dually, groisser1997instantons, ciaglia2021focus, murray1993information, matsueda2017banados, naudts2009q, petz2007quantum, lesniewski1999monotone, facchi2010classical, barbaresco2014koszul]).
Information geometry exhibits deep connections with several mathematical disciplines, including Hessian geometry, Kähler geometry, complex geometry, special Kähler geometry, and complex analysis. One of the most fundamental problems in information geometry is the realization of a statistical manifold as a statistical model, that is, a manifold equipped with dual connections with respect to a Fisher metric. This question was resolved by Hông Vân Lê [le2006statistical]. Another, much more challenging problem is to determine which classes of statistical manifolds can be realized as exponential families; this problem remains open. Despite its rapid development, information geometry still contains many significant open problems [sei2025open].
This paper is organized as follows. Section 2 reviews the relevant background notions in information geometry. Section 3 is devoted to the proof of the main result of the paper.
2. Basics Notions
2.1. Statistical models
Definition 2.1.
[amari1987differential, amari2000methods, amari2012differential, amari2014curvature, amari2016information, ay2017information, calin2014geometric] Let $\mathcal{S}$ be a family of probability distributions. Assume that $\mathcal{S}$ is represented as
$\mathcal{S}=\{p(x,\theta)\mid\theta\in\Theta\},$
where $\Theta$ is a domain in $\mathbb{R}^{n}$ with Cartesian coordinates $(\theta^{1},\theta^{2},\dots,\theta^{n})$ (called the parameter space). In addition, we assume that the mapping $\theta\mapsto p(\cdot,\theta)$ is bijective. Then $\mathcal{S}$ is said to be a statistical model of probability distributions determined by $p$.
The Fisher information metric $g^{F}$ on $\mathcal{S}$ is defined in local coordinates $\theta=(\theta^{1},\dots,\theta^{n})$ by
$g^{F}_{ij}(\theta)=\mathbb{E}_{p(\cdot,\theta)}\left[\frac{\partial\log p(x,\theta)}{\partial\theta^{i}}\,\frac{\partial\log p(x,\theta)}{\partial\theta^{j}}\right],$
where $\mathbb{E}_{p(\cdot,\theta)}$ denotes expectation with respect to $p(\cdot,\theta)$. We assume:
  1. (i)
    $g^{F}_{ij}(\theta)$ is finite and smooth,
  2. (ii)
    the matrix $\big(g^{F}_{ij}(\theta)\big)$ is positive definite for all $\theta\in\Theta$.
The corresponding Riemannian metric is
$g^{F}=\sum_{i,j=1}^{n}g^{F}_{ij}(\theta)\,d\theta^{i}\otimes d\theta^{j},$
defines a Riemannian metric on $\mathcal{S}$, called the Fisher metric of $\mathcal{S}$.
For $\alpha\in\mathbb{R}$, the $\alpha$-connection $\nabla^{(\alpha)}$ on a statistical model $\mathcal{S}$ is defined by
$\Gamma^{(\alpha)}_{ijk}(\theta)=\mathbb{E}_{p(\cdot,\theta)}\left[\left(\frac{\partial^{2}\log p(x,\theta)}{\partial\theta^{i}\partial\theta^{j}}+\frac{1-\alpha}{2}\frac{\partial\log p(x,\theta)}{\partial\theta^{i}}\frac{\partial\log p(x,\theta)}{\partial\theta^{j}}\right)\log p(x,\theta)\cdot\frac{\partial\log p(x,\theta)}{\partial\theta^{k}}\right].$
In the literature, the family of $\alpha$-connections $\nabla^{(\alpha)}$ is commonly referred to as the Amari-Chentsov $\alpha$-connections or Amari–Čencov $\alpha$-connections [amari2000methods] . This terminology honors the foundational contributions of Shun-ichi Amari [amari1987differential, amari2012differential] and Nikolai Chentsov[chentsov1982statistical] to information geometry. The Amari-Chentsov $\alpha$-connections provide a unified framework for studying the differential geometric structure of statistical models, with the Fisher metric as the central Riemannian structure and the $\alpha$-connections forming a dualistic structure on the statistical manifold.
The $\alpha$-connection $\nabla^{(\alpha)}$ satisfies the following fundamental properties:
  1. (1)
    $\nabla^{(\alpha)}$ is torsion-free for all $\alpha\in\mathbb{R}$.
  2. (2)
    $\nabla^{(0)}$ coincides with the Levi-Civita connection of the Fisher metric.
Among all $\alpha$-connections, the $\pm 1$-connections play a particularly important role in the geometric theory of statistical inference. $\nabla^{(1)}$ is called the exponential connection (or e-connection). $\nabla^{(-1)}$ is called the mixture connection (or m-connection).
It’s well known that $\alpha$-connections, the following duality formula holds:
$\partial_{\theta_{k}}g^{F}(\partial_{\theta_{i}},\partial_{\theta_{j}})=g^{F}\!\left(\nabla^{(\alpha)}_{\partial_{\theta_{k}}}\partial_{\theta_{i}},\,\partial_{\theta_{j}}\right)\;+\;g^{F}\!\left(\partial_{\theta_{i}},\,\nabla^{(-\alpha)}_{\partial_{\theta_{k}}}\partial_{\theta_{j}}\right).$
The connections $\nabla^{(\alpha)}$ and $\nabla^{(-\alpha)}$ are said to be dual (or conjugate) with respect to $g^{F}$.
For any $\alpha,\gamma\in\mathbb{R}$, the Christoffel symbols of the Amari-Chentsov connections $\nabla^{(\alpha)}$ and $\nabla^{(\gamma)}$ are related by:
$\Gamma^{(\gamma)}_{ij,k}=\Gamma^{(\alpha)}_{ij,k}+\frac{\alpha-\gamma}{2}\,\text{T}^{F}_{ijk},$
where
$\text{T}^{F}_{ijk}(\theta)=\mathbb{E}_{p(\cdot,\theta)}\left[\partial_{i}\ell_{\theta}\,\partial_{j}\ell_{\theta}\,\partial_{k}\ell_{\theta}\right],\quad\text{with }\ell_{\theta}(x)=\log p(x,\theta).$
The fully symmetric $(0,3)$-tensor field $\text{T}^{F}$ defined by the components $\text{T}^{F}_{ijk}$ is called the cubic form or Amari-Chentsov tensor. Moreover, the covariant derivative of the Fisher metric with respect to $\nabla^{(\alpha)}$ satisfies:
$(\nabla^{(\alpha)}_{X}g^{F})(Y,Z)=\alpha\,\text{T}^{F}(X,Y,Z).$
An important and widely studied class of statistical models is the exponential family, consisting of probability densities of the form
$p(x,\theta)=\exp\left[C(x)+\sum_{i=1}^{n}\theta^{i}F_{i}(x)-\psi(\theta)\right],\quad\theta\in\Theta,$
where $C(x),F_{1}(x),\dots,F_{n}(x)$ are smooth functions of the observation $x$, and $\psi$ is the cumulant generating function on the parameter space $\Theta$. The coordinates $\theta=(\theta^{1},\dots,\theta^{n})$ are called natural parameters, and $\psi$ is determined by the normalization condition:
$\psi(\theta)=\log\int\exp\left[C(x)+\sum_{i=1}^{n}\theta^{i}F_{i}(x)\right]dx.$
A particularly important special case is the normal distribution, obtained by taking
$C(x)=0,\quad F_{1}(x)=x,\quad F_{2}(x)=x^{2},\quad\theta^{1}=\frac{\mu}{\sigma^{2}},\quad\theta^{2}=-\frac{1}{2\sigma^{2}},$
which yields
$\psi(\theta^{1},\theta^{2})=-\frac{(\theta^{1})^{2}}{4\theta^{2}}+\frac{1}{2}\log\left(-\frac{\pi}{\theta^{2}}\right).$
This corresponds exactly to the Gaussian distribution $\mathcal{N}(\mu,\sigma)$.
For an exponential family with natural parameters $\theta=(\theta^{1},\dots,\theta^{n})$, the Fisher metric and the cubic form are given in terms of the cumulant generating function $\psi(\theta)$ by:
$g^{F}_{ij}(\theta)=\partial_{i}\partial_{j}\psi(\theta),\quad\text{T}^{F}_{ijk}(\theta)=\partial_{i}\partial_{j}\partial_{k}\psi(\theta),$
where $\partial_{i}=\frac{\partial}{\partial\theta^{i}}$.
2.2. Statistical manifold
Building upon the geometric insights of Amari and Chentsov, Steffen L. Lauritzen provided a formalize the notion of a statistical manifold.
Definition 2.2 (Lauritzen [lauritzen1987statistical]).
A statistical manifold is a quadruple $(M,g,\nabla,\nabla^{*})$ where:
  1. (1)
    For all $X,Y,Z\in\mathfrak{X}(M)$,
    $X(g(Y,Z))=g(\nabla_{X}Y,Z)+g(Y,\nabla^{*}_{X}Z).$
  2. (2)
    Both $\nabla$ and $\nabla^{*}$ are torsion-free affine connections.
Equivalently, a statistical manifold may be defined as a triplet $(M,g,\nabla)$ such that $\text{T}=\nabla g$ is a totally symmetric $(0,3)$-tensor field.
For a statistical manifold $(M,g,\nabla,\nabla^{*})$, the triple $(g,\nabla,\nabla^{*})$ is called a statistical structure, and the pair $(\nabla,\nabla^{*})$ are referred to as statistical connections. The Levi–Civita connection $\nabla^{\mathrm{LC}}$ satisfies
(2.1) $2\,\nabla^{\mathrm{LC}}=\nabla+\nabla^{*}.$
Note that the $(0,3)$-tensor field T is referred to by Lauritzen [lauritzen1987statistical] as the skewness tensor.
The following construction provides a fundamental example of how statistical structures naturally arise on Riemannian manifolds equipped with a distinguished vector field.
Let $(M,g)$ be a Riemannian manifold with Levi–Civita connection $\nabla^{\mathrm{LC}}$, and let $\xi\in TM\setminus\{0\}$ be a non-vanishing vector field. For all $X,Y\in\mathfrak{X}(M)$, define
(2.2) $\nabla_{X}Y$ $=\nabla^{\mathrm{LC}}_{X}Y+g(X,\xi)g(Y,\xi)\,\xi,$
(2.3) $\nabla^{*}_{X}Y$ $=\nabla^{\mathrm{LC}}_{X}Y-g(X,\xi)g(Y,\xi)\,\xi.$
Then $(M,g,\nabla,\nabla^{*})$ is a statistical manifold.
Starting from our previous construction, on an almost contact metric manifold $(M^{2d+1},\phi,g,\alpha,\xi)$ with $\nabla^{LC}$ the Levi–Civita connection of the metric $g$ (see [blair2010riemannian] for exposition on almost contact metric geometry), we can construct the following one-parameter families of statistical connections:
(2.4) $\nabla^{\varepsilon}$ $=\nabla^{LC}+\varepsilon\,\alpha(\cdot)\alpha(\cdot)\xi,\qquad\forall\varepsilon\in\mathbb{R},$
(2.5) $\nabla^{-\varepsilon}$ $=\nabla^{LC}-\varepsilon\,\alpha(\cdot)\alpha(\cdot)\xi,\qquad\forall\varepsilon\in\mathbb{R}.$
For a statistical manifold $(M,g,\nabla,\nabla^{*})$. The connection $\nabla$ is flat if and only if $\nabla^{*}$ is flat. In this case, we say that $(M,g,\nabla,\nabla^{*})$ is a dually flat manifold ([amari2016information]).
It was shown in [ay2002dually] that any compact dually flat manifold necessarily has an infinite fundamental group. Consequently, the sphere $\mathbb{S}^{q}$, for $q>1$, cannot be a dually flat manifold.
2.3. Hessian manifold
Hessian manifolds exhibit particularly rich topological and geometric structures and share a deep relationship with Kähler manifolds. This relationship is especially manifested in the fact that the tangent bundle of any Hessian manifold is naturally endowed with a Kähler metric induced by the Hessian structure itself [shima2007geometry, dombrowski1962geometry, satoh2019almost]. Such manifolds occupy a central place in information geometry, Kähler geometry, special Kähler geometry, and complex analysis. They also play an important role in various applied fields, including convex optimization, statistics, combinatorics, Souriau’s thermodynamics on Lie groups, and quantitative finance.
Definition 2.3.
A locally flat manifold is a pair $(M,\nabla)$, where $\nabla$ is a connection whose torsion and curvature tensors vanish identically:
$T^{\nabla}=0,\qquad R^{\nabla}=0.$
Such a connection $\nabla$ is called a locally flat connection, and $(M,\nabla)$ is referred to as a locally flat manifold.
Definition 2.4.
An affinely flat structure on an $m$-dimensional manifold $M$ is a complete atlas
(2.6) $\mathcal{A}=\{(U_{i},\Phi_{i})\},$
such that the transition maps $\Phi_{i}^{-1}\circ\Phi_{j}$ belong to the group $\text{Aff}(\mathbb{R}^{n})$ of affine transformations on $\mathbb{R}^{m}$.
It’s important to note that a manifold admits an affine structure if and only if it is locally flat.
Definition 2.5 ([shima1997geometry, shima2007geometry]).
A Riemannian metric $g$ on a locally flat manifold $(M,\nabla)$ is called a Hessian metric if, locally, there exists a smooth function $f$ such that
$g=\nabla^{2}f,\quad\text{i.e.,}\quad g_{ij}=\frac{\partial^{2}f}{\partial x_{i}\partial x_{j}},$
where $(x_{1},\dots,x_{n})$ is a system of affine coordinates with respect to $\nabla$.
A locally flat manifold $(M,\nabla)$ equipped with a Hessian metric $g$ is called a Hessian manifold and is denoted by $(M,g,\nabla)$.
Let $D$ denote the Levi-Civita connection of $(M,g)$, and define
$\gamma=D-\nabla.$
Since both $\nabla$ and $D$ are torsion-free, we have
$\gamma_{X}Y=\gamma_{Y}X,\quad\text{for all vector fields }X,Y.$
Moreover, in affine coordinates, the components $\gamma^{i}_{jk}$ of $\gamma$ coincide with the Christoffel symbols $\Gamma^{i}_{jk}$ of $D$.
Proposition 2.1 ([shima1997geometry]).
Let $(M,\nabla)$ be an locally flat manifold and $g$ a Riemannian metric. The following are equivalent:
  1. (1)
    $g$ is a Hessian metric;
  2. (2)
    $(\nabla_{X}g)(Y,Z)=(\nabla_{Y}g)(X,Z)$ for all vector fields $X,Y,Z$;
  3. (3)
    In affine coordinates $(x^{i})$, $g_{ij}$ satisfies
    $\frac{\partial g_{ij}}{\partial x^{k}}=\frac{\partial g_{kj}}{\partial x^{i}};$
  4. (4)
    $g(\gamma_{X}Y,Z)=g(Y,\gamma_{X}Z)$ for all vector fields $X,Y,Z$;
  5. (5)
    $\gamma_{ijk}=\gamma_{jik}$.
On a orientable Hessian manifold $(M,g,\nabla)$, using the connection $\nabla$ and the volume element induced by $g$, Koszul introduced a closed 1-form, which Shima [shima2007geometry] refers to as the first Koszul form, associated with the Hessian structure $(\nabla,g)$.
Definition 2.6 ([koszul1961domaines, shima1997geometry]).
The first Koszul form $\beta$ associated with $(\nabla,g)$ is the $1$-form defined by
$\nabla_{X}(\mathrm{Vol}_{g})=\beta(X)\,\mathrm{Vol}_{g},\quad X\in\mathfrak{X}(M),$
where $\mathrm{Vol}_{g}$ denotes the Riemannian volume form induced by $g$.
Directly from the definition,
$\beta(X)=\operatorname{Tr}(\gamma_{X}),$
and in affine coordinates,
$\beta_{i}=\frac{1}{2}\frac{\partial\log\det(g_{pq})}{\partial x_{i}}=\gamma^{k}_{ki}.$
We are now able to present the main theorem.
3. Main Results
Obviously, a statistical model is a statistical manifold. Lauritzen and Amari raised the following fundamental question: can every statistical manifold be realized as a finite-dimensional statistical model ? Constructing a realization of a statistical manifold as a statistical model, i.e., a manifold with dual connections with respect to a Fisher metric, is an important question in information geometry. While a positive answer to this problem was given by the work of Hông Vân Lê [le2006statistical], explicitly constructing the probability family that gives rise to the Fisher metric remains a challenging task in general.
Theorem 3.1 ([le2006statistical]).
Every smooth compact statistical manifold $(M^{n},g,T)$ of finite dimension admits, for a sufficiently large integer $N$, an isostatistical embedding (i.e., an embedding that preserves the tensor $T$) into the space $\mathrm{Cap}_{N}$ of probability densities on $\Omega_{N}$, where $\Omega$ is a differentiable manifold endowed with a Borel measure.
A major challenge in information geometry is to identify which classes of statistical manifolds $(M,g,T)$ arise from an exponential family. To date, to the best of our knowledge, this problem remains unsolved.
We can now state the main result of this paper.
Theorem 3.2.
On any compact statistical manifold, there exists a singular foliation whose leaves are Hessian manifolds. Every compact, orientable leaf is either a finite quotient of a flat torus (flat) or a mapping torus with periodic glueing map. If such a leaf is three-dimensional and non-flat, then all of its Betti numbers are odd, and it arises as a quotient of an exponential family.
Proof.
The proof proceeds in four steps.
  1. Step 1:
    Foliations by Hessian manifolds.
    Let $(M,g,\nabla)$ be a statistical manifold. The equation associated with $\nabla$ is defined by:
    $\mathrm{E}(\nabla):\quad\nabla^{2}X=0\quad\text{for all }X\in\mathfrak{X}(M),$
    where $\nabla^{2}$ denotes the second covariant derivative, given for all $X,Y,Z\in\mathfrak{X}(M)$ by:
    $\nabla^{2}_{X,Y}Z:=\nabla_{X}(\nabla_{Y}Z)-\nabla_{\nabla_{X}Y}Z.$
    Let $(x^{1},\dots,x^{m})$ be local coordinates on $M$ and $X\in\mathfrak{X}(M)$. Setting
    $\partial_{i}$ $=\frac{\partial}{\partial x^{i}},$ $X$ $=\sum_{k}X^{k}\partial_{k},$ $\nabla_{\partial_{i}}\partial_{j}$ $=\sum_{k}\Gamma^{k}_{ij}\partial_{k},$
    the princip al symbol of the operator $X\mapsto\nabla^{2}X$ takes the form:
    $(\nabla^{2}X)(\partial_{i},\partial_{j})=\sum_{k}\theta^{k}_{ij}(X)\partial_{k},$
    where
    $\theta^{l}_{ij}(X)$ $=\frac{\partial^{2}X^{l}}{\partial x^{i}\partial x^{j}}+\Gamma^{l}_{ik}\frac{\partial X^{k}}{\partial x^{j}}+\Gamma^{l}_{jk}\frac{\partial X^{k}}{\partial x^{i}}-\Gamma^{k}_{ij}\frac{\partial X^{l}}{\partial x^{k}}$
    $\quad+\frac{\partial\Gamma^{l}_{jk}}{\partial x^{i}}+\Gamma^{m}_{jk}\Gamma^{l}_{im}-\Gamma^{m}_{ij}\Gamma^{l}_{mk}.$
    The equation $\mathrm{E}(\nabla)$ thus reduces locally to the system $\theta^{l}_{ij}(X)=0$.
    Let $\mathscr{J}_{\nabla}(M)$ denote the sheaf of solutions to the equation $\mathrm{E}(\nabla)$ and $J_{\nabla}=\Gamma(M,\mathscr{J}_{\nabla})$ its space of global sections. Define a product on $J_{\nabla}$ by:
    $X\cdot Y=\nabla_{X}Y.$
    A direct computation shows:
    $X\cdot(Y\cdot W)-(X\cdot Y)\cdot W=\nabla_{X}(\nabla_{Y}W)-\nabla_{\nabla_{X}Y}W=\nabla^{2}_{X,Y}W.$
    Hence, for all $W\in J_{\nabla}$,
    $X\cdot(Y\cdot W)-(X\cdot Y)\cdot W=0.$
    Now, for $U,V\in J_{\nabla}$, we have:
    $X\cdot(Y\cdot(U\cdot V))-(X\cdot Y)\cdot(U\cdot V)$ $=X\cdot((Y\cdot U)\cdot V)-((X\cdot Y)\cdot U)\cdot V$
    $=[X\cdot(Y\cdot U)-(X\cdot Y)\cdot U]\cdot V=0.$
    Moreover, $\nabla_{U}V\in J_{\nabla}$, since:
    $\nabla^{2}_{X,Y}(\nabla_{U}V)=X\cdot(Y\cdot(U\cdot V))-(X\cdot Y)\cdot(U\cdot V)=0.$
    Thus, $(J_{\nabla},\cdot)$ is an associative algebra. The principal symbol analysis shows that $\mathrm{E}(\nabla)$ is of finite type [guillemin1965integrability, singer1965infinite, kumpera1972lie], so $J_{\nabla}$ is finite-dimensional. As $\nabla$ is torsion-free, the bracket
    $X,Y]_{\nabla}=\nabla_{X}Y-\nabla_{Y}X$
    coincides with the Lie bracket, and $J_{\nabla}$ is closed under it, forming a finite-dimensional Lie subalgebra of $\mathfrak{X}(M)$. By Lie’s third theorem, there exists a unique simply connected Lie group $G_{\nabla}$ with Lie algebra isomorphic to $J_{\nabla}$. Since $M$ is compact, $J_{\nabla}$ is integrable [palais1957global], yielding a locally effective differentiable action:
    $G_{\nabla}\times M\to M.$
    Let $\mathcal{F}_{\nabla}$ be the singular foliation defined by the orbits of $G_{\nabla}$, and $g_{\nabla}$ the restriction of $g$ to the leaves of $\mathcal{F}_{\nabla}$. Since $J_{\nabla}$ is an associative algebra with respect to $\nabla$, each leaf is $\nabla$-autoparallel. Let $\nabla^{\mathcal{F}_{\nabla}}$ denote the restriction of $\nabla$ to the leaves. A straightforward computation shows that $\nabla^{\mathcal{F}_{\nabla}}_{X}g_{\nabla}$ is totally symmetric. Moreover, for $U\in J_{\nabla}$, we have
    $R^{\nabla^{\mathcal{F}_{\nabla}}}(X,Y)U=\nabla^{\mathcal{F}_{\nabla},2}_{X,Y}U-\nabla^{\mathcal{F}_{\nabla},2}_{Y,X}U=0\quad\text{for all }U\in J_{\nabla},$
    where
    $\nabla^{\mathcal{F}_{\nabla},2}_{X,Y}Z=\nabla^{\mathcal{F}_{\nabla}}_{X}\big(\nabla^{\mathcal{F}_{\nabla}}_{Y}Z\big)-\nabla^{\mathcal{F}_{\nabla}}_{\nabla^{\mathcal{F}_{\nabla}}_{X}Y}Z.$
    Hence, the leaves of $\mathcal{F}_{\nabla}$ are dually flat manifolds, and thus Hessian manifolds.
    Then, on any statistical manifold $(M,g,\nabla)$, there always exists a singular foliation $\mathcal{F}_{\nabla}$ whose leaves are Hessian manifolds. A similar construction is available for the dual connection $\nabla^{*}$. This completes the first step of the proof of the theorem.
  2. Step 2:
    Topology of compact orientable leaves.
    Let $L_{\nabla}$ be a compact orientable leaf of the singular foliation $\mathcal{F}_{\nabla}$. By Step 1, the triple $(L_{\nabla},g_{\nabla},\nabla^{\mathcal{F}_{\nabla}})$ is a compact orientable Hessian manifold. Let $D$ denote the Levi-Civita connection of $g_{\nabla}$. The first Koszul form $\beta\in\Omega^{1}(L_{\nabla})$ associated with the statistical structure $(\nabla^{\mathcal{F}_{\nabla}},g_{\nabla})$ is defined by the relation
    $\nabla^{\mathcal{F}_{\nabla}}_{X}(\mathrm{Vol}_{g_{\nabla}})=\beta(X)\,\mathrm{Vol}_{g_{\nabla}},\quad\text{for all }X\in\mathfrak{X}(L_{\nabla}),$
    where $\mathrm{Vol}_{g_{\nabla}}$ is the Riemannian volume form induced by $g_{\nabla}$ [koszul1961domaines, shima1997geometry].
    According to Shima and Yagi [shima1996geometry, Theorem 4.1], the first Koszul form $\beta$ is $D$-parallel, i.e., $D\beta=0$. Consequently, its norm $\|\beta\|_{g_{\nabla}}$ is constant on $L_{\nabla}$. This leads to two mutually exclusive cases:
    Case 1: Vanishing Koszul form ($\beta=0$).
    By [shima1996geometry, Theorem 4.2], the identity $\beta=0$ implies that the Levi-Civita connection coincides with the flat affine connection: $D=\nabla^{\mathcal{F}_{\nabla}}$. Hence $(L_{\nabla},g_{\nabla})$ is a flat Riemannian manifold. Since $L_{\nabla}$ is compact, the classical theorems of Bieberbach [bieberbach1911bewegungsgruppen, bieberbach1912bewegungsgruppen] and Wolf [Wolf] imply that $L_{\nabla}$ is isometric to a flat torus:
    $L_{\nabla}\simeq\mathbb{R}^{d}/\Lambda,$
    where $d=\dim L_{\nabla}$ and $\Lambda\subset\mathbb{R}^{d}$ is a Bieberbach group, acting freely on $\mathbb{R}^{d}$ and $\simeq$ , means “isometric to.” In other words, $L_{\nabla}$ is finitely covered by a flat torus, such as $\mathbb{T}^{d}=\mathbb{R}^{d}/\mathbb{Z}^{d}$.
    Case 2: Non-vanishing Koszul form ($\beta\neq 0$).
    In this case, $\beta$ is a non-vanishing $D$-parallel 11-form. Define the vector field $Z$ dual to $\beta$ via the metric: $\beta=g_{\nabla}(Z,\cdot)$. Since $D\beta=0$, it follows that $DZ=0$, i.e., $Z$ is a parallel vector field on $L_{\nabla}$.
    By Welsh’s classification [welsh1986manifolds, Theorem 1], the existence of a non-trivial parallel vector field on a compact Riemannian manifold implies that $L_{\nabla}$ is a fiber bundle over $\mathbb{S}^{1}$ with finite structure group denoted $G$. Furthermore, the results of Tischler [tischler1970fibering] imply that $L_{\nabla}$ is diffeomorphic to a mapping torus:
    $L_{\nabla}\cong X_{\varphi}=\frac{X\times[0,1]}{(p,0)\sim(\varphi(p),1)},$
    with the associated fibre bundle
    $X\rightarrow L_{\nabla}=X_{\varphi}\rightarrow\mathbb{S}^{1},$
    where $\varphi\in\mathrm{Diff}(X)$. By [bazzoni2014structure, Proposition 6.4], the structural group $G$ generates a finite cyclic subgroup $\langle\varphi\rangle\subset\mathrm{Diff}(X)$. From [welsh1986manifolds], we deduce that $\varphi$ is of finite order, and in particular periodic. This concludes the second step.
  3. Step 3:
    Three-dimensional non-flat leaves.
    Now assume that the leaf $L_{\nabla}$ is 3-dimensional and non-flat. In this case, the first Koszul form $\beta$ is a non-vanishing closed 1-form. Define the corresponding non-vanishing vector field $Z$ by $\beta=g_{\nabla}(Z,\cdot)$. Since $D\beta=0$, it follows that $DZ=0$, and consequently $g_{\nabla}(Z,Z)=c>0$ is constant.
    Define the normalized vector field and 1-form by:
    $\widetilde{Z}:=\frac{Z}{\sqrt{c}},\qquad\widetilde{\beta}:=\frac{\beta}{\sqrt{c}}.$
    Then we have:
    $\widetilde{\beta}(\widetilde{Z})=1,\quad D\widetilde{Z}=0,\quad\mathcal{L}_{\widetilde{Z}}g_{\nabla}=0,\quad g_{\nabla}(\widetilde{Z},\widetilde{Z})=1.$
    Define a 2-form by:
    $\Omega=i_{\widetilde{Z}}\mathrm{Vol}_{g_{\nabla}},$
    where $\mathrm{Vol}_{g_{\nabla}}$ denotes the Riemannian volume form. A straightforward computation shows:
    $d\Omega=d(i_{\widetilde{Z}}\mathrm{Vol}_{g_{\nabla}})=\mathcal{L}_{\widetilde{Z}}\mathrm{Vol}_{g_{\nabla}}=0.$
    Observe that:
    $i_{\widetilde{Z}}(\widetilde{\beta}\wedge\mathrm{Vol}_{g_{\nabla}})=0,$
    which immediately yields:
    $\mathrm{Vol}_{g_{\nabla}}=\widetilde{\beta}\wedge\Omega.$
    We thus deduce that $\widetilde{Z}$ is the Reeb vector field associated with the cosymplectic structure $(\widetilde{\beta},\Omega)$ (see [li2008topology] ). Since $\widetilde{Z}$ is a Killing vector field on $L_{\nabla}$, it follows from [bazzoni2015k] that $(L_{\nabla},\widetilde{\beta},\Omega)$ is a K-cosymplectic manifold. By [bazzoni2015k, Proposition 2.8], $L_{\nabla}$ admits an almost contact metric structure $(\widetilde{\beta},\widetilde{Z},\phi,h)$ with $\mathcal{L}_{\widetilde{Z}}h=0$.
    From Goldberg [goldberg1969integrability, Proposition 3], the structure $(\widetilde{\beta},\widetilde{Z},\phi,h)$ is normal (i.e., $N_{\phi}=0$); hence $(L_{\nabla},\widetilde{\beta},\widetilde{Z},\phi,h)$ is a co-Kähler manifold ( Kähler-mapping torus [li2008topology]). Furthermore, by [Chinea1993TopologyOC, Theorem 11], the first Betti number $b_{1}(L_{\nabla})$ is odd. Applying Poincaré duality, it follows that all Betti numbers $b_{s}(L_{\nabla})$ are odd.
    From [cornu1987varietes] (Proposition 2), $(L_{\nabla},\widetilde{\beta},\Omega)$ cannot be a contact manifold. Then from [geiges1997normal] (Theorems 1 and 2), we deduce that $L_{\nabla}$ is diffeomorphic to one of the following:
    1. (a)
      $\Lambda\backslash(\mathbb{H}^{2}\times\mathbb{R})$ with $\Lambda\subset\mathrm{Isom}_{0}(\mathbb{H}^{2}\times\mathbb{R})$,
    2. (b)
      $\mathbb{T}^{2}$-bundles over $\mathbb{S}^{1}$ with periodic monodromy,
    3. (c)
      $\mathbb{S}^{2}\times\mathbb{S}^{1}$.
    Since $(L_{\nabla},g_{\nabla},\nabla^{\mathcal{F}_{\nabla}})$ is a Hessian manifold, Shima [shima1981hessian] shows that $\mathbb{S}^{2}\times\mathbb{S}^{1}$ does not admit a Hessian metric, so this case is excluded.
    Now assume $L_{\nabla}$ is a $T^{2}$-bundle over $\mathbb{S}^{1}$ with periodic monodromy. From [martelli2016introduction] (Exercise 11.4.13), we deduce that $L_{\nabla}=\mathbb{T}^{2}_{A}$, where $A\in\mathrm{SL}_{2}(\mathbb{Z})$ with $A=\pm I$ or $|\mathrm{tr}(A)|<2$. Then from [martelli2016introduction] (Proposition 11.4.14), $L_{\nabla}$ is a Seifert manifold with Euler number zero and whose base orbifold has zero Euler characteristic. By [martelli2016introduction] (Theorem 12.3.1), $L_{\nabla}$ admits a flat metric. Since $L_{\nabla}$ is non-flat by assumption, this case is also excluded.
    Therefore, the only remaining possibility is:
    $L_{\nabla}\cong\Gamma\backslash(\mathbb{H}^{2}\times\mathbb{R})\quad\text{with}\quad\Gamma\subset\mathrm{Isom}_{0}(\mathbb{H}^{2}\times\mathbb{R})\cong\mathrm{Isom}_{0}(\mathbb{H}^{2})\times\mathrm{Isom}_{0}(\mathbb{R}).$
  4. Step 4:
    Exponential family structure.
    Following [molnar1997projective], we identify the manifold $\mathbb{H}^{2}\times\mathbb{R}$ with the Lorentz cone ( the positive light-cone)
    $\Theta=\left\{(x,y,z)\in\mathbb{R}^{3}:z>0,\ z^{2}>x^{2}+y^{2}\right\}.$
    By [shima2007geometry, Example 4.2], $\Theta$ is a regular self-dual convex cone. Then for $\theta=(x,y,z)\in\Theta$ and $\theta^{*}=(\xi,\eta,\zeta)\in\Theta$, consider the exponential family on $\Theta$ defined by
    $P((x,y,z),(\xi,\eta,\zeta))=\frac{e^{-(x\xi+y\eta+z\zeta)}}{\chi(x,y,z)},$
    where $\chi$ denotes the Koszul-Vinberg characteristic function ([vinberg1967theory, shima2007geometry, koszul1968deformations, koszul1961domaines, koszul1962ouverts, barbaresco2014koszul, faraut1994analysis, satake2014algebraic, vesentini1976invariant]):
    $\chi(x,y,z)=\iiint\limits_{\begin{subarray}{c}\zeta^{2}>\xi^{2}+\eta^{2}\\[2.0pt] \zeta>0\end{subarray}}\exp\!\left[-(\xi x+\eta y+\zeta z)\right]\,d\xi\,d\eta\,d\zeta.$
    Let $q=z^{2}-x^{2}-y^{2}$. From [shima2007geometry, faraut1994analysis], we obtain the explicit expression
    $\chi(x,y,z)=q^{-3/2}\chi(0,0,1),$
    which yields the probability density
    $P((x,y,z),(\xi,\eta,\zeta))=\exp\left(-(x\xi+y\eta+z\zeta)+\frac{3}{2}\log q-\log\chi(0,0,1)\right).$
    Thus, $\{P(\theta,\theta^{*}):\theta\in\Theta\}$ forms an exponential family parametrized by $\Theta$. Its Fisher metric is given by
    $g^{F}_{ij}(\theta)=\mathbb{E}\left[\frac{\partial\log P}{\partial\theta^{i}}\frac{\partial\log P}{\partial\theta^{j}}\right],\quad\theta\in\Theta.$
    A direct computation shows that
    $g^{F}=-\frac{3}{2}\,(\nabla^{0})^{2}\log q,$
    where $\nabla^{0}$ is the Levi–Civita connection of $\mathbb{R}^{3}$. Hence, $(\Theta,\nabla^{0},g^{F})$ is a Hessian manifold and $g^{F}$ is a Koszul Hessian metric. In coordinates, the metric is
    $\begin{split}g^{F}=\frac{3}{q^{2}}\Big[&(z^{2}+x^{2}+y^{2})\,dz^{2}-4zx\,dz\,dx-4zy\,dz\,dy\\ &+(z^{2}+x^{2}-y^{2})\,dx^{2}+4xy\,dx\,dy+(z^{2}-x^{2}+y^{2})\,dy^{2}\Big].\end{split}$
    Introducing cylindrical coordinates $(t,r,\alpha)$ by
    $\begin{cases}z=e^{t}\cosh r,\\ x=e^{t}\sinh r\cos\alpha,\\ y=e^{t}\sinh r\sin\alpha,\end{cases}\quad t\in\mathbb{R},\ r\geq 0,\ \alpha\in[0,2\pi),$
    the metric simplifies to
    $g=3(dt^{2}+dr^{2}+\sinh^{2}r\,d\alpha^{2}).$
    Here $(r,\alpha)$ are polar coordinates on $\mathbb{H}^{2}$ and $t$ is the fiber coordinate. Thus, $(\mathbb{H}^{2}\times\mathbb{R},g)$ is isometric to $(\Theta,g^{F})$.
    We deduce that, the leaf $L_{\nabla}$ of the associated foliation $\mathcal{F}_{\nabla}$ satisfies
    $L_{\nabla}\cong\Lambda\backslash\Theta,\qquad\Lambda\subset\mathrm{Isom}_{0}(\Theta).$
    Identifying $\theta\in\Theta$ with the probability distribution $P_{\theta}$, we obtain
[…]
    where $\{P_{\theta}:\theta\in\Theta\}$ is an exponential family. This completes the proof.
Remark 3.3.
For the reader familiar with the geometries of $3$-manifolds in the sense of Thurston, the leaf
$L_{\nabla}\cong\Gamma\backslash(\mathbb{H}^{2}\times\mathbb{R}),\qquad\Gamma\subset\mathrm{Isom}_{0}(\mathbb{H}^{2}\times\mathbb{R}),$
is a Seifert manifold with Euler number 0 and hyperbolic orbifold base $B$ (since $\chi_{\mathrm{orb}}(B)<0$). In other words, $L_{\nabla}$ admits a $\mathbb{H}^{2}\times\mathbb{R}$-geometry.
From [martelli2016introduction, Prop. 10.3.26], $L_{\nabla}$ is covered by $\Sigma_{g}\times\mathbb{S}^{1}$ for some $g>1$. Hence
$L_{\nabla}\cong\Pi\backslash(\Sigma_{g}\times\mathbb{S}^{1}),$
where $\Pi$ is a finite group acting freely and properly on $\Sigma_{g}\times\mathbb{S}^{1}$.
Since $\Sigma_{g}$ admits an $\mathbb{RP}^{2}$-structure (see [benzecri1960varietes, p. 243] and [goldman2022geometric, Cor. 6.5.2]), the universal cover of $\Sigma_{g}\times\mathbb{S}^{1}$, namely $\mathbb{H}^{2}\times\mathbb{R}$, develops into the Lorentz cone $\Theta$. Consequently,
$\Sigma_{g}\times\mathbb{S}^{1}\cong(\Gamma\times\langle\lambda\rangle)\backslash\Theta,\qquad\Gamma\subset\mathrm{SO}(2,1),\quad\lambda>1.$
We deduce that
$L_{\nabla}\cong\Pi\backslash\bigl((\Gamma\times\langle\lambda\rangle)\backslash\Theta\bigr).$
Finally, by using the Koszul–Vinberg characteristic function $\chi$ on the sharp convex cone $\Theta$, and arguing as above, the result follows.